# Matryoshka Query Transformer for Large Vision-Language Models

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) typically encode an image into a fixed number of visual tokens (e.g., 576) and process these tokens with a language model. Despite their strong performance, LVLMs face challenges in adapting to varying computational constraints. This raises the question: can we achieve flexibility in the number of visual tokens to suit different tasks and computational resources? We answer this with an emphatic yes. Inspired by Matryoshka Representation Learning, we introduce the Matryoshka Query Transformer (MQT), capable of encoding an image into $m$ visual tokens during inference, where $m$ can be any number up to a predefined maximum. This is achieved by employing a query transformer with $M$ latent query tokens to compress the visual embeddings. During each training step, we randomly select $m\leq M$ latent query tokens and train the model using only these first $m$ tokens, discarding the rest. Combining MQT with LLaVA, we train a single model once, and flexibly and drastically reduce the number of inference-time visual tokens while maintaining similar or better performance compared to training independent models for each number of tokens. Our model, MQT-LLaVA, matches LLaVA-1.5 performance across 11 benchmarks using a maximum of 256 tokens instead of LLaVA’s fixed 576. Reducing to 16 tokens (8x less TFLOPs) only sacrifices the performance by 2.4 points on MMBench. On certain tasks such as ScienceQA and MMMU, we can even go down to only 2 visual tokens with performance drops of just 3% and 6% each. Our exploration of the trade-off between the accuracy and computational cost brought about by the number of visual tokens facilitates future research to achieve the best of both worlds. Our code and model are publicly available  [^1] https://github.com/gordonhu608/MQT-LLaVA
@@ 1 Introduction
Recent work in Large Vision-Language Models (LVLMs) (OpenAI, 2023; Liu et al., 2023b; Bai et al., 2023) has shown remarkable performance across a broad range of vision-language tasks (Huang et al., 2023; Chen et al., 2023; Cai et al., 2024; Li et al., 2023b). These LVLMs typically consist of a vision encoder to embed images into grid features, which are fed into a Large Language Model (LLM) (Touvron et al., 2023; Chiang et al., 2023) for processing and reasoning alongside a text input.
@@ 5.1 Large Vision Language Model
Large Vision Language Models (LVLMs), pioneered by Liu et al. (2023b); Zhu et al. (2023); Dai et al. (2023) have successfully showcased promising results on a wide variety of vision-language perception and reasoning tasks. Recent works further expand the capabilities of LVLMs to region-level image understanding (Huang et al., 2023; Peng et al., 2023; Chen et al., 2023; Cai et al., 2024), video understanding  (Zhang et al., 2023; Li et al., 2023b; Jin et al., 2024) and 3D understanding (Hong et al., 2023; Szot et al., 2024). These models mostly consist of a vision encoder and an LLM aligned by a vision-language connector module, which can be an MLP (Liu et al., 2023b; Hu et al., 2023), Q-Former or queries through cross attentions (Dai et al., 2023; Bai et al., 2023; Ye et al., 2023), or Resampler (Alayrac et al., 2022). The number of visual tokens output by these modules can be very large, especially for higher image resolutions, multiple frames in video tasks, and multiple images for in-context learning. In this paper, we propose a training paradigm that can dynamically choose a number of visual tokens that adapts to variable computation costs at inference time.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) typically encode an image into a fixed number of visual tokens (e.g., 576) and process these tokens with a language model. Despite their strong performance, LVLMs face challenges in adapting to varying computational constraints. This raises the question: can we achieve flexibility in the number of visual tokens to suit different tasks and computational resources? We answer this with an emphatic yes. Inspired by Matryoshka Representation Learning, we introduce the Matryoshka Query Transformer (MQT), capable of encoding an image into $m$ visual tokens during inference, where $m$ can be any number up to a predefined maximum. This is achieved by employing a query transformer with $M$ latent query tokens to compress the visual embeddings. During each training step, we randomly select $m\leq M$ latent query tokens and train the model using only these first $m$ tokens, discarding the rest. Combining MQT with LLaVA, we train a single model once, and flexibly and drastically reduce the number of inference-time visual tokens while maintaining similar or better performance compared to training independent models for each number of tokens. Our model, MQT-LLaVA, matches LLaVA-1.5 performance across 11 benchmarks using a maximum of 256 tokens instead of LLaVA’s fixed 576. Reducing to 16 tokens (8x less TFLOPs) only sacrifices the performance by 2.4 points on MMBench. On certain tasks such as ScienceQA and MMMU, we can even go down to only 2 visual tokens with performance drops of just 3% and 6% each. Our exploration of the trade-off between the accuracy and computational cost brought about by the number of visual tokens facilitates future research to achieve the best of both worlds.
@@ 1 Introduction
Recent work in Large Vision-Language Models (LVLMs) (OpenAI, 2023; Liu et al., 2023b; Bai et al., 2023) has shown remarkable performance across a broad range of vision-language tasks (Huang et al., 2023; Chen et al., 2023; Cai et al., 2024a; Li et al., 2023b). These LVLMs typically consist of a vision encoder to embed images into grid features, which are fed into a Large Language Model (LLM) (Touvron et al., 2023; Chiang et al., 2023) for processing and reasoning alongside a text input.
@@ 5.1 Large Vision Language Model
Large Vision Language Models (LVLMs), pioneered by Liu et al. (2023b); Zhu et al. (2023); Dai et al. (2023); Yin et al. (2023) have successfully showcased promising results on a wide variety of vision-language perception and reasoning tasks. Recent works further expand the capabilities of LVLMs to region-level image understanding (Huang et al., 2023; Peng et al., 2023; Chen et al., 2023; Cai et al., 2024a), video understanding  (Zhang et al., 2023; Li et al., 2023b; Jin et al., 2024) and 3D understanding (Hong et al., 2023; Szot et al., 2024). These models mostly consist of a vision encoder and an LLM aligned by a vision-language connector module, which can be an MLP (Liu et al., 2023b; Hu et al., 2023), Q-Former or queries through cross attentions (Dai et al., 2023; Bai et al., 2023; Ye et al., 2023), or Resampler (Alayrac et al., 2022). The number of visual tokens output by these modules can be very large, especially for higher image resolutions, multiple frames in video tasks, and multiple images for in-context learning. In this paper, we propose a training paradigm that can dynamically choose a number of visual tokens that adapts to variable computation costs at inference time.
@@ 5.2 Efficient Vision Transformers
Concurrent work (Cai et al., 2024b) also studies matryoshka-style visual tokens, and designs 5 scales of pooling layers to control the granularity of images. Different from their work, we introduce a query transformer to extract visual tokens, enabling a more flexible choice of any number of visual tokens under a predefined maximum. Their work corroborates our findings by demonstrating robust performance and efficient use of a minimal number of visual tokens.